assume the scores on an achievent test taken by fifth graders are normally distributed with M=57 an Sd=7.5.
a. what % of student scored above the mean/
b.above what raw score do 5% of scores fall?
c. what % of students scored below a score of 50?
d. between what 2 raw scores do the middle 50% of these scores fall.
need to undersatnd the correct way to solve these types of problems.
green_m1@yahoo.com

Answer by stanbon(75887)

assume the scores on an achievent test taken by fifth graders are normally distributed with M=57 an Sd=7.5.
-----------------------------------------------------------
a. what % of student scored above the mean:: you always have 50% above the mean.
---------------------------
b. above what raw score do 5% of scores fall?
Assuming you mean the top 5%:
invNorm(0.95) = 1.645
score = 1.756*7.5+57 = 69.34
---------------------------
c. what % of students scored below a score of 50?
z(50) = (50-57)/7.5 = -7/7.5 = -0.9333
P(x < 50 = P(z < -0.9333) = normalcdf(-100,-0.9333) = 0.1753
------------------------
d. between what 2 raw scores do the middle 50% of these scores fall.
Lower limit has a left tail of 25%
The corresponding z-value is invnorm(0.25) = -0.6745
Lower limit score = -0.6745*7.5+57 = 51.94
-------------------
Upper limit has a right tail of 25%
Upper limit score = +0.6745*7.5+57 = 62.06
-----------------------------
Cheers,
Stan H.
need to undersatnd the correct way to solve these types of problems.